xXMarcelieXx:

i need some decimal help how did they get 23.21 billion ? ( 320.7 million ) ( 72.36 ) which i got 23205.852

Shadow:

You did \[320.7 \times 72.36 \] ?

xXMarcelieXx:

yes which the 320.7 was million

Shadow:

How did you write that out

xXMarcelieXx:

i just multiplied it as (320.7) (72.36)

Shadow:

Okay, you got 23205.852 because you did not properly represent 320.7 million. You can represent it with:\[320.7 \times 10^6\]

xXMarcelieXx:

oh okay. which is 320,700,000

Shadow:

Then multiply it by 72.36

xXMarcelieXx:

then i got 2.3205852E10 so then we move one decimal to right?

Shadow:

I got 23,205,852,000

xXMarcelieXx:

oh hmmm how did they get the decimal

Shadow:

You move the decimal 10 times to the right.

xXMarcelieXx:

oml lol

xXMarcelieXx:

can you show me how you would move them

xXMarcelieXx:

@Shadow com back

Shadow:

2.3205852000 

xXMarcelieXx:

oh okay but what happen to the decimal there?

Shadow:

What do you mean?

xXMarcelieXx:

im confuse why the answer is in this decimal form 23.21

Shadow:

23.21 billion is the same as: \[23.21 \times 10^9 = 23, 210, 000,000\]

Shadow:

It's just a different way of saying it.

xXMarcelieXx:

ohh so its just the 2 numbers from the beginning

22west:

yeah